OpenStudy (anonymous):

HELP ME! Does anyone know a simple way of solving systems involving circles?

hero (hero):

Post your question and we'll see

OpenStudy (anonymous):

x^2+y^2=36 y=2x+15

hero (hero):

So I guess substitution is not the easy way?

hero (hero):

x^2 + (2x+15)^2 = 36

OpenStudy (anonymous):

Well it is but once it comes to factoring it takes a little longer then i thought

hero (hero):

Well there is another way. Let's see if it works.

OpenStudy (anonymous):

I got 5x^2+60x+189. Now i have to factor but it isn't coming out right. And i am trying the quadratic formula.

hero (hero):

square the second equation to get y^2 = (2x+15)^2 y^2 = 36 - x^2 y^2 = (2x+15)(2x+15) y^2 = 2x(2x+15)+15(2x+15) y^2 = 4x^2 + 30x + 30x + 225 y^2 = 4x^2 + 60x + 225 y^2 = y^2 4x^2 + 60x + 225 = 36 - x^2 5x^2 + 60x + 189 = 0 ac = 189*5 a + c = 60 That leads to a dead end unfortunately

hero (hero):

I'm convinced that the roots are complex

OpenStudy (anonymous):

They are and I dont know to go around that so i am lost.

hero (hero):

There's no way around complex roots.

OpenStudy (anonymous):

True. But for the answer Do I just put zero?

hero (hero):

The answer to this depends. What exactly is the question that you were required to answer. Tell me word for word.

OpenStudy (anonymous):

All it says is solve the systems and it just have a list of different problems for me to do.

hero (hero):

Have you been studying complex roots at all?

hero (hero):

If it doesn't say, you'll have to assume it wants the complex roots

OpenStudy (anonymous):

Yes I master that. Lol. But for solving systems involving circles is very complex to understand.

hero (hero):

If you can find the roots like you say you can, what exactly is the difficulty?

hero (hero):

Maybe it just annoys you

OpenStudy (anonymous):

Because I am questioning Do my teacher wants complex numbers part of my answer.

OpenStudy (anonymous):

It kind is because i been stuck on this for a while now.

hero (hero):

I'm pretty sure that's what your teacher wants. You have no choice but to provide them since he or she didn't specify.

OpenStudy (anonymous):

True so that means my answer will be complex for the quadratic formula.

hero (hero):

In this case, you have to provide the (x,y) points

OpenStudy (anonymous):

Yes because that part of my solution.

hero (hero):

There's no easy way to calculate the complex roots. As far as I'm concerned, you have three options: 1. Quad Formula 2. Complete The Square 3. Myininaya's Method Trust me. You don't want Myininaya's Method.

OpenStudy (anonymous):

Ok. Then thanks for your help if i still need help i just hop back online.

hero (hero):

lol; it said "false" on my calculator. I should have investigated further why it said that.

OpenStudy (anonymous):

@telliott99 I drew it out also and i got that. Now what would be my (x,y)

OpenStudy (anonymous):

I must confess, that I haven't read all the posts here with equal intensity, but if I combine what @Hero and @telliott99 suggested already, it seems like these two functions never meet. One is a line, the other is a circle. The solution is that there is no solution, therefore the line never intersects the circle.

OpenStudy (anonymous):

So in case your textbook says different @TammyFlow, you might want to check the question again and see if there wasn't any confusion of \(\pm \) signs.

OpenStudy (anonymous):

I got that by the picture i drew. so thanks @Spacelimbus . And to @TheViper I know how to do that already but thanks for the help

OpenStudy (theviper):

yw dude :)

OpenStudy (anonymous):

It wasn't a textbook. It was just a sheet with problems on there about solving systems involving circles . So i am just going to draw it out for now on

OpenStudy (anonymous):

@TheViper I am a girl not a dude.

OpenStudy (theviper):

LOL so wht should I say I don't know English so much :(

OpenStudy (anonymous):

@TheViper Well you can your welcome instead of putting dude.

OpenStudy (theviper):

OK u r welcome :)

OpenStudy (theviper):

@TammyFlow :D

OpenStudy (anonymous):

Ah okay, basically, you can do the drawing out version, but algebraically works just fine. A circle with a circle always ends up in a quadratic equation. Therefore compute the discriminant: \[ \Large D= b^2-4ac\] If D=0 you have found a tangent to the circle, it only meets the circle at exactly one point If D>0 then you can compute the quadratic equation normally and you will obtain two solutions, therefore you get two points where your line meets the circle, it transverses it. Finally if D<0 then there are only complex roots, which means that there are no solutions in \( \mathbb{R}\) and your line never meets the circle at any given point.

OpenStudy (anonymous):

@TheViper ;). See that's better

OpenStudy (anonymous):

So i guess ill try that way also. But for some problems ill just draw it out

OpenStudy (anonymous):

whatever suits you the best (-:

OpenStudy (anonymous):

Thanks for your help. I'll see which one is simple.

hero (hero):

Or you can just use a graphing calculator and graph both equations. Graphing is the fastest way to see whether the line and the circle actually intersect. The reason why I didn't graph it first is because I expected the two equations to have a solution. When the solutions were complex, for some reason it didn't occur to me it was because the line and the circle did not intersect. Must've been smoking something yesterday.

OpenStudy (anonymous):

No it was just late. Now i know graphing is easier then anything depending on the equation. But thanks for all your help.

OpenStudy (anonymous):

@TammyFlow It's always useful to sketch. It may not be necessary, but it keeps your feet on the ground. And, FWIW :) you're not a dude, your're a dudette. At least, that's what we say in my family..

OpenStudy (anonymous):

@telliott99 Thanks Dude now i have earn a new nickname. And it is more useful to graph it because it save time depending on the equation.

OpenStudy (anonymous):

Hope it's ok. Graphing is really the first thing to do almost always.

OpenStudy (anonymous):

Ok. Thanks